# Pose Anything: A Graph-Based Approach for Category-Agnostic Pose Estimation

### v1

@@ Abstract
Traditional 2D pose estimation models are limited by their category-specific design, making them suitable only for predefined object categories. This restriction becomes particularly challenging when dealing with novel objects due to the lack of relevant training data.
To address this limitation, category-agnostic pose estimation (CAPE) was introduced. CAPE aims to enable keypoint localization for arbitrary object categories using a single model, requiring minimal support images with annotated keypoints. This approach not only enables object pose generation based on arbitrary keypoint definitions but also significantly reduces the associated costs, paving the way for versatile and adaptable pose estimation applications.
We present a novel approach to CAPE that leverages the inherent geometrical relations between keypoints through a newly designed Graph Transformer Decoder. By capturing and incorporating this crucial structural information, our method enhances the accuracy of keypoint localization, marking a significant departure from conventional CAPE techniques that treat keypoints as isolated entities.
We validate our approach on the MP-100 benchmark, a comprehensive dataset comprising over 20,000 images spanning more than 100 categories. Our method outperforms the prior state-of-the-art by substantial margins, achieving remarkable improvements of 2.16% and 1.82% under 1-shot and 5-shot settings, respectively. Furthermore, our method’s end-to-end training demonstrates both scalability and efficiency compared to previous CAPE approaches.
1 Introduction
@@
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Pose Anything. Given a support image and skeleton from any category (purple) our model localizes the skeleton on a query image (green). Our graph-oriented category-agnostic pose estimation method integrates structural information to improve keypoint localization. Model was trained on one group of object classes and tested on different classes. Test set results are shown here.
@@
Current CAPE methods match support keypoints with their corresponding counterparts in the query image. The matching is done in latent space to facilitate the localization of keypoints. One notable distinction from previous CAPE methods lies in their treatment of keypoints as individual, disconnected points. In contrast to these approaches, we recognize the significance of the underlying geometrical structure of objects. This structure serves as a robust prior that can significantly enhance the accuracy of keypoint localization. Building upon this realization, we introduce a novel Graph Transformer Decoder, purposefully designed to capture and incorporate this crucial structural information. By doing so, we aim to exploit the inherent relationships and dependencies between keypoints.
@@
Our method outperforms the previous best approach CapeFormer [32] by large margins, with an improvement of 2.16% and 1.82% under 1-shot and 5-shot settings, respectively. Our method is trained end-to-end and demonstrates both scalability and efficiency compared with previous approaches.
@@
  • We introduced a new method, leveraging the geometrical relations between keypoints using a novel Graph Transformer Decoder.
@@
  • State-of-the-art performance on the MP-100 benchmark.
@@ 2.1 Detection Transformer
The DEtection TRansformer (DETR) [3] stands out for its simplicity and conceptually straightforward and universally applicable approach, requiring minimal domain-specific knowledge, and avoiding the need for customized label assignments or non-maximum suppression.
However, the original DETR design suffers from challenges such as slow convergence rates and reduced detection accuracy. In response, a multitude of subsequent studies, including [24, 11, 7, 53, 50, 20], have emerged, improving the DETR paradigm. These works have paved the way for the development of top-tier object detectors, leveraging innovations like the reintroduction of multi-scale features and local cross-attention computations [50, 9, 39].
A compelling connection emerges between detection problems, focused on bounding box estimation and localization, and pose estimation, which centers on keypoint localization. Recognizing this connection, we incorporate several advancements from the DETR model into the latest category-agnostic pose estimation method [32], thus improving its performance.
@@ 2.2 Category-Agnostic Pose Estimation
The primary goal of pose estimation is to accurately localize the semantic keypoints of objects or instances. Traditionally, pose estimation methods have predominantly been tailored to specific categories, such as humans [8, 2, 46], animals [49, 48], or vehicles [34, 29]. Existing research focused mainly on designing powerful convolutional neural networks [8, 2, 16] or transformer-based architectures [23, 43]. However, these methods are limited to object categories encountered during training.
A relatively unexplored facet is category-agnostic pose estimation, introduced by POMNet [42], where the emphasis shifts towards constructing a generic representation and similarity metrics. This approach predicts keypoints by comparing support keypoints with query images in the embedding space, addressing the challenge of object categories not seen during training.
@@
POMNet [42] employs a transformer to encode query images and support keypoints, with a regression head predicting similarity directly from the concatenation of support keypoint and query image features. CapeFormer [33] extends the matching paradigm to a two-stage framework, rectifying unreliable matching outcomes to enhance prediction precision. We follow CapeFormer, enhancing their methodology with a more robust baseline. Moreover, we focus on the importance of geometrical structure, seamlessly integrating it into our architectural framework.
2.3 Graph Neural Network
@@
Originally designed for graph data like social networks [12], citation networks [30], and biochemical graphs [37], GCNs are gaining traction in computer vision tasks [41, 17, 38, 15]. They find a variety of applications such as scene graph generation, point cloud classification, and action recognition. Scene graph generation involves parsing images into graphs that represent objects and their relations, often combined with object detection [41, 45]. Point clouds, derived from LiDAR scans, are effectively classified and segmented using GCNs [17, 40]. Furthermore, GCNs have been employed for human activity recognition by processing naturally constructed graphs of human joints [44, 5, 21]. Recent work even suggests using GNNs as backbone feature extractors for visual tasks, representing images as graph structures [13].
In contrast to GCNs, which uniformly aggregate neighbor information based on the adjacency matrix, Graph Attention Networks (GAT)[36] introduce the concept of self-attention. GAT incorporates the graph structure into the attention mechanism through masked attention. Similarly, aGCN[45] relies on self-attention to determine the weights of neighboring joints, with differences arising from their choice of activation functions and transformation matrices. While GAT and aGCN advance the aggregation of neighbor joints, the challenge of limited receptive fields persists. Zhaoet al. [51] addressed this limitation by introducing a learnable weighting matrix to the adjacency matrix, enabling the learning of semantic relationships between 2D joints. Similarly, Lin et al.[19] leveraged a standard transformer encoder by making adjustments to the dimension of encoder layers. However, these approaches, as previously mentioned, fail to fully exploit the inherent graph structure. In an attempt to address this challenge, Graformer [52] suggested a solution involving a combination of attention blocks followed by graph convolutions that employ learnable adjacency matrices. Nevertheless, their approach is primarily tailored for domain-specific applications, making it less suitable for our category-agnostic task.
Our category-agnostic method tackles this challenge by integrating both self-attention and a GCN network into the decoder’s architecture. Self-attention promotes learnable connections between keypoints, while the GCN, using the input from the skeleton graph, aids the model in sharing information among keypoints that are semantically related.
@@ 3 Method
We start with a short description of an enhanced baseline that we build on top of CapeFormer [32]. Then, we describe a graph-based approach that takes advantage of the strong graph structure available in the data. The complete architecture of our method is illustrated in Figure 2.
3.1 An Enhanced Baseline
We make two changes to CapeFormer architecture. First, we replace its Resnet-50 [14] backbone with a comparable-sized stronger, transformer-based, SwinV2-T [22] backbone. We also remove positional encoding because we found that it introduces an undesirable dependency on the order of the keypoints. Using these modifications, we create an Enhanced Baseline of the previous CapeFormer architecture. We elaborate on the impact of these changes in the experimental section.
For completeness, we give a brief description of CapeFormer, whos design is similar to the DETR architecture and consists of 4 sub-networks:
  • Shared backbone. A pre-trained Resnet-50 is used to extract features from the input support and query images. To acquire the support keypoint features, we perform an element-wise multiplication between the support image feature map and keypoint masks. These masks are created by positioning Gaussian kernels centered at the support keypoints locations. In scenarios involving multiple support images, such as 5-shot settings, we calculate the average (in feature space) of the support keypoint features from different images. This results in both the query feature map $\hat{F_{q}}\in\mathbb{R}^{hw\times C}$ and the support keypoint features $\hat{F_{s}}\in\mathbb{R}^{K\times C}$.
  • Transformer encoder. A transformer encoder is used to fuse information between the support keypoint features and the query patch features. The encoder consists of three transformer blocks, each housing a self-attention layer. Before entering the self-attention layer, the input support keypoints and query features are concatenated and subsequently separated once more. The outputs are refined query feature map $F_{q}$ and refined support keypoint features $F_{s}$.
  • Similarity-Aware Proposal Generator. The proposal generator aligns support keypoint features with query features, resulting in similarity maps. From these maps, we select peaks to serve as the similarity-aware proposals. In pursuit of a balance between efficiency and versatility, a trainable inner-product mechanism [31] is employed to explicitly model similarity.
  • Transformer decoder. To decode keypoint locations from the query feature map, a transformer decoder network is utilized. The Transformer decoder consists of three layers, with each layer including self-attention, cross-attention, and feed-forward blocks. An iterative refinement strategy, inspired by prior work [1, 35, 53], is applied to enable each decoder layer to predict coordinate deltas for the previous coordinate predictions. Additionally, similar to the approach in Conditional DETR [25], the decoder utilizes the predicted coordinates to provide enhanced reference points for pooling features from the image feature map.
@@ 3.2 A Graph-Based Approach
The core idea of our work is to take advantage of the geometrical structure encoded in the pose graph. Our method is built upon the enhanced baseline, replacing the transformer decoder module with our novel Graph Transformer Decoder. We now discuss our new Graph Transformer Decoder and the associated loss function.
Graph Transformer Decoder
Figure 3: Graph Transformer Decoder. GTD is based on the original CapeFormer decoder, changing the feed-forward network from a simple MLP to a graph-based network. (a) A scheme of the transformer decoder which includes self-attention, cross-attention, and a feed-forward network. Self-attention encourages adaptive interactions among support keypoints, while cross-attention extracts localization information. (b) Previous feed-forward consisted of an MLP with 2 layers. (c) Our novel feed-forward includes a GCN layer and subsequent linear layers that enhance keypoint features and promote information exchange among known connected keypoints.
We introduce a Graph Transformer Decoder (GTD) that works directly on the graph structure. As can be seen in Figure 3, GTD utilizes a novel feed-forward network in the transformer decoder layer, that includes a Graph Convolutional Network Layer (GCN).
We recognize that self-attention, a mechanism that helps our model focus on relevant information, can be thought of as a graph convolutional network (GCN) with a learnable adjacency matrix. When we are dealing with pose estimation for a single category, this mechanism is sufficient for learning the relationships between keypoints and integrating a learned structure into the model. However, for category-agnostic pose estimation (CAPE) tasks, where the model needs to work with various object categories, it is beneficial to explicitly consider the semantic connections between keypoints. This helps the model break symmetry, maintain consistent structure, and handle noisy keypoints by sharing information among neighboring keypoints.
@@
We implemented this prior in the transformer decoder. Specifically, GTD is based on the original CapeFormer decoder, changing the feed-forward network from a simple MLP to a GCN network. To address the potential problem of excessive smoothing often observed in deep GCNs [18, 27], which can lead to a reduction in the distinctiveness of node characteristics and consequently a decline in performance, we introduce a linear layer for each node following the GCN layer. Figure 4 illustrates the power of implementing the graph prior in the cross-attention layers. As can be seen, early layers focus on a keypoint, while later layers take into consideration also neighboring keypoints (according to the graph structure).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) (b) (c) (d) (e)
Support Support 1st Layer Last Layer Predicted
Skeleton Keypoint Heatmap Heatmap Location
Figure 4: Attention Map Visualization. We illustrate the cross-attention within the decoder, specifically between $F^{\prime}_{s}$ and $F_{q}$. (a) illustrated the support skeleton, and (b) visualizes the searched keypoint in the support image. (c) In the initial layer, attention is centered solely on the queried keypoints. (d) In contrast, the final layer expands attention to additional keypoints based on skeletal connectivity. (e) Our decoder adeptly utilizes structural cues to accurately position keypoints within the query image.
@@
The original decoder has three main components: self-attention, cross-attention, and a feed-forward network. Self-attention allows for adaptive interactions among support keypoints, transforming the input feature $F_{s}^{l}$ into $F_{s}^{\prime}$. Cross-attention augments support keypoints’ representation by extracting localization information from query feature patches, enhancing attention at proposed locations. This involves concatenating the keypoint localization embedding with keypoint features. Cross-attention inputs are $F_{s}^{\prime}$ (with the localization embedding) as queries, patch features $F_{q}$ (with positional encoding) as keys, and $F_{q}$ as values, resulting in transformed features $F_{s}^{\prime\prime}$. Lastly, we employ a novel feed-forward network to process the output keypoint features. To incorporate prior geometrical knowledge, we introduce a graph neural network (GCN) layer. This GCN layer further concentrates the keypoint features, facilitating the exchange of information between known keypoint connections.
@@
Where $W_{i}\in\mathbb{R}^{C_{out}\times C_{in}}$ is a learnable parameter matrix, $\sigma_{act}$ is an activation function (ReLU), and $\widetilde{A}\in\mathbb{R}^{K\times K}$ is the symmetrically normalized form of the adjacency matrix $A\in[0,1]^{K\times K}$, which is defined as $a_{ij}=1$ for node $v_{j}$ that is connected to node $v_{i}$ and 00 elsewhere.
The output features $F_{s}^{l+1}$ are used to update the keypoints locations from $P^{l}$ to $P^{l+1}$. We follow [32] and use the following update function:
@@
Training loss
Following CapeFormer, we use two supervision signals: a heatmap loss and an offset loss. The heatmap loss supervises the similarity metric and initial coordinates suggestion, while the offset loss supervises the localization output:

### v2

# https://orhir.github.io/pose-anything/

@@ Abstract
Traditional 2D pose estimation models are limited by their category-specific design, making them suitable only for predefined object categories. This restriction becomes particularly challenging when dealing with novel objects due to the lack of relevant training data. To address this limitation, category-agnostic pose estimation (CAPE) was introduced. CAPE aims to enable keypoint localization for arbitrary object categories using a few-shot single model, requiring minimal support images with annotated keypoints.
We present a significant departure from conventional CAPE techniques, which treat keypoints as isolated entities, by treating the input pose data as a graph. We leverage the inherent geometrical relations between keypoints through a graph-based network to break symmetry, preserve structure, and better handle occlusions. We validate our approach on the MP-100 benchmark, a comprehensive dataset comprising over 20,000 images spanning over 100 categories. Our solution boosts performance by 0.98% under a 1-shot setting, achieving a new state-of-the-art for CAPE. Additionally, we enhance the dataset with skeleton annotations. Our code and data are publicly available.
@@
Figure 1: Given a support image and skeleton from any category (purple) our model localizes the skeleton on a query image (green). Our graph-oriented method integrates structural information to improve keypoint localization.
Keywords:
Class-Agnostic Pose Estimation, Few-Shot Learning
1 Introduction
@@
CAPE methods match support keypoints with their corresponding counterparts in the query image. The matching is done in latent space to facilitate the localization of keypoints. Previous CAPE methods treat keypoints as individual, disconnected points. However, inherent geometrical relations between keypoints serve as a robust prior that can enhance the accuracy of keypoint localization by breaking symmetry and handling occlusions. In contrast to these approaches, we recognize this significant geometrical structure and leverage it by treating the input keypoints as a graph. Building upon this realization, we introduce GraphCape, a graph-based model, purposefully designed to capture and incorporate this crucial structural information. By doing so, we exploit the inherent relationships and dependencies between keypoints. We explored various design choices, detailed in the supplementary, and ultimately chose to incorporate this prior into the model architecture.
@@
It is important to note that the required additional keypoints connectivity input is shared among instances of the same category. Thus, the additional graph annotations required equals the number of categories. For example, when using our method for fine-tuning or inference on a specific category, the annotation process requires adding only one skeletal definition, which is quite simple.
We compare our method to previous CAPE methods, as well as an updated version of the previous state-of-the-art which we call CapeFormer-T. Our method improves over this updated version, boosting performance under a 1-shot setting by 0.98%, achieving a new state-of-the-art.
@@
  • We propose to treat the input keypoints as connected nodes of a graph, instead of independent entities. Furthermore, to leverage the graph connectivity we introduce a new graph-based model design.
@@
  • We achieve state-of-the-art performance on the MP-100 benchmark.
@@ 2.1 Detection Transformer
A compelling connection emerges between detection problems, focused on bounding box estimation and localization, and pose estimation, which centers on keypoint localization. Recognizing this connection, we incorporate several advancements from object detection models into the latest category-agnostic pose estimation method [29], thus improving its performance.
The DEtection TRansformer (DETR)[2] was the first transformer-based network for object detection, replacing conventional spatial anchors with learnable object queries in a transformer decoder. Its simplicity and universally applicable approach require minimal domain-specific knowledge, avoiding customized label assignments or non-maximum suppression. However, the original DETR design faced challenges such as slow convergence rates and reduced detection accuracy. In response, numerous studies[24, 10, 6, 48, 45, 20] have emerged to improve the DETR paradigm. These works have led to the development of top-tier object detectors, leveraging innovations like the reintroduction of multi-scale features and local cross-attention computations [45, 8, 34].
@@ 2.2 Category-Agnostic Pose Estimation
The primary goal of pose estimation is to localize the semantic keypoints of objects or instances accurately. Traditionally, pose estimation methods have predominantly been tailored to specific categories, such as humans [7, 1, 41], animals [44, 43], or vehicles [31, 27]. Existing research focused mainly on designing convolutional neural networks [7, 1, 15] or transformer-based architectures [23, 38]. However, these methods are limited to categories encountered during training.
A relatively unexplored area is category-agnostic pose estimation (CAPE), introduced by Xu et al. [37]. This approach predicts keypoints by comparing support keypoints with query images in the embedding space, handling unseen object categories during training. This shifts pose estimation into few-shot learning, which utilizes prototype, meta-learning, and fine-tuning. ProtoNet [30] learns a prototype for each class in the support data and classifies query data based on the nearest prototype. MAML [9] is a meta-learning-based approach that finds optimal initialization weights for rapid generalization to novel tasks with minimal fine-tuning. Another few-shot approach is fine-tuning [25], where the model is pre-trained on all base categories and fine-tuned on support images of novel categories during testing.
@@
Specifically designed for CAPE, POMNet [37] employs a transformer to encode query images and support keypoints, with a regression head predicting similarity directly from the concatenation of support keypoint and query image features. CapeFormer [29] extends the matching paradigm to a two-stage framework, rectifying unreliable matching outcomes to enhance prediction precision. We first update CapeFormer, enhancing their methodology. Then, we focus on the importance of geometrical structure, integrating it into our model design.
2.3 Graph Neural Networks in Computer Vision
@@
Originally designed for graph data like social networks [11], citation networks [28], and biochemical graphs [32], GCNs have become pivotal in computer vision tasks [36, 16, 33, 14], finding applications in scene graph generation, point cloud classification, and action recognition. Scene graph generation parses images into graphs representing objects and their relations, often combined with object detection [36, 40]. Point clouds from LiDAR scans are effectively classified and segmented using GCNs [16, 35]. Additionally, GCNs process human joints’ graphs for activity recognition [39, 4, 22]. Recent work even suggests using GNNs as backbone feature extractors for visual tasks, representing images as graphs [12].
Graph-driven models have also been proposed for 6D pose estimation to better infuse structure information [3, 46, 47]. They utilize orientation-aware autoencoders with 3D graph convolutions, which are robust to point shift and object size variations and construct k-NN graphs, extracting geometry-aware inter-modality correlations. Lian et al. [18] adopts GNNs to model interactions among 3D keypoints, aiming for 2D correspondences in input images.
Inspired by these approaches, we integrate GCNs into the decoder’s architecture to enhance semantic connections between keypoints and share information among structurally close ones.
@@ 3 Method
In the following section, we describe our graph-based approach that takes advantage of the strong graph structure available in the data. The complete architecture of our method is illustrated in Figure 2. Our approach utilizes a pre-trained backbone to extract support keypoints and query image features, followed by a transformer encoder that refines these features through self-attention. A similarity proposal generator is employed for initial coordinates localization. Finally, a transformer decoder predicts the keypoints’ locations. More details on the different components are in the supplementary.
@@ 3.1 A Graph-Based Approach
The core idea of our work is in treating the input keypoints as a graph, to take advantage of the geometrical structure encoded in the graph connectivity. Our key insight is in recognizing that self-attention, a mechanism that helps models focus on relevant information, can be thought of as a GCN with an input-dependent learnable adjacency matrix. When we are dealing with pose estimation for a single category, this mechanism is sufficient for learning the relationships between keypoints and integrating a learned structure into the model. However, for the new task of CAPE, where the model needs to work with object categories it has never seen before, it is beneficial to explicitly consider the semantic connections between keypoints. As GNN promotes information sharing between connected keypoints, location data is shared among structurally connected keypoints, which are typically close. Thus, the graph design aids in structure preservation and occlusion handling. Moreover, we believe that utilizing the graph structure during training helps in breaking feature symmetry, as it enforces similar features based on structure (e.g., right/left) rather than solely on semantics. Figure 3 illustrates this hypothesis qualitatively. Notably, our graph-based method, despite being trained on multiple categories, exhibits similar attention patterns to models trained on single categories.
Refer to caption Refer to caption Refer to caption
Multi-Category Single-Category Multi-Category
Base Model Base Model Graph Model
Figure 3: Self-Attention Map Visualization. Comparing self-attention in decoders of three models: (a) CapeFormer-T trained on various object categories, (b) CapeFormer-T trained only on furniture objects, and (c) GraphCape trained on various object categories using a graph structure. Observe the edges between the legs and base of the chair. Notably, our graph-based method, despite being trained on multiple categories, exhibits similar attention patterns to models trained on single categories.
@@
We implemented this prior into the transformer decoder. Specifically, our design is based on CapeFormer, changing the decoder’s feed-forward network from a simple MLP to a GCN-based module. To address the potential problem of excessive smoothing often observed in deep GCNs [17, 26], which can lead to a reduction in the distinctiveness of node characteristics and consequently a decline in performance, we introduce a linear layer for each node following the GCN layer. Figure 4 illustrates our new design.
Figure 4: Graph FFN. The Transformer decoder is based on the original CapeFormer design, changing the feed-forward network from a simple MLP to a graph-based network. (a) A scheme of the transformer decoder which includes self-attention, cross-attention, and a feed-forward network. Self-attention encourages adaptive interactions among support keypoints, while cross-attention extracts localization information. (b) Previous FFN consisted of an MLP with 2 layers. (c) Our graph FFN includes a GCN layer and subsequent linear layers that enhance keypoint features and promote information exchange among known connected keypoints.
@@
The decoder has three main components: self-attention, cross-attention, and a feed-forward network. Self-attention allows for adaptive interactions among support keypoints, transforming the input feature $F_{s}^{l}$ into $F_{s}^{\prime}$. Cross-attention augments support keypoints’ representation by extracting localization information from query feature patches, enhancing attention at proposed locations. This involves concatenating the keypoint localization embedding with keypoint features. Cross-attention inputs are $F_{s}^{\prime}$ (with the localization embedding) as queries, patch features $F_{q}$ (with positional encoding) as keys, and $F_{q}$ as values, resulting in transformed features $F_{s}^{\prime\prime}$. Lastly, we employ a graph-based feed-forward network to process the output keypoint features as a graph. This layer further concentrates the keypoint features, facilitating the exchange of information between neighboring keypoints in the graph.
@@
Where $W_{i}\in\mathbb{R}^{C_{out}\times C_{in}}$ is a learnable parameter matrix, $\sigma_{act}$ is an activation function (ReLU), and $\widetilde{A}\in\mathbb{R}^{K\times K}$ is the symmetrically normalized form of the adjacency matrix $A\in[0,1]^{K\times K}$. $A$ is a binary matrix, defined as $a_{ij}=1$ for node $v_{j}$ that is connected to node $v_{i}$ and 00 elsewhere. While the adjacency matrix is fixed for each category, during training some nodes may not be visible in the support image. Thus, we also mask the adjacency matrix before normalization, to make sure only meaningful information is propagated between nodes.
The output features $F_{s}^{l+1}$ are used to update the normalized keypoints locations from $P^{l}$ to $P^{l+1}$. We follow [29] and use the following update function:
@@
3.2 Training Scheme
Following CapeFormer, we use two supervision signals: a heatmap loss and an offset loss. The heatmap loss supervises the proposal generator, constraining the shape of the similarity maps, and facilitating the learning of meaningful representations, while the offset loss supervises the localization output:
@@ 3.2 Training Scheme
3.3 Encoding Structure: Positional Encoding VS Graph Network
Table 1: Keypoint Positional Encoding: Comparison of results using CapeFormer with the original dataset and using permuted keypoints. Keypoint Positional Encoding (PE) adds an implicit structure bias according to the fixed order of keypoints in the dataset. Graph-FFN adds an explicit structure prior of any object’s shape.
Original Data Shuffled Keypoints
With PE 89.26 64.20 (-25.06%)
Without PE 85.24 85.24 (0%)
Without PE + Graph-FFN 87.71 87.71 (0%)
CapeFormer introduced a keypoint positional encoding (PE), termed the "Support Keypoint Identifier". This encoding is generated via sinusoidal encoding of the keypoints’ order. We analyzed this feature thoroughly and present the results in Table 1. As shown, this encoding improves performance by a significant 4% PCK (defined later), at the cost of introducing a dependency on the order of keypoints. This became evident when we evaluated a model trained with this positional encoding on data of keypoints in permuted order. Using PE and shuffling the keypoints results in a notable decrease of approximately 25% in PCK, indicating that the model had become highly specialized for a specific keypoint format. This sharp decrease is not present when the positional encoding is not used. Although the dataset contains diverse object categories, some of them share a common keypoint ordering. This indicates that using this PE adds structure information, learned from known categories in the training set.
As we require invariance to keypoint permutations, graph structure is an obvious choice. Our graph-based approach serves as an alternative, boosting performance while being agnostic to permutations in the order of nodes.